Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
SPB
SQS
XO
ToBeMigrated2XWiki
ToBeMigrated2GitlabPages
ToBeMigratedByGroup
Migrated2Zulip
Archived2PDF
not2MigrateOrArchive
SCS
SXP Migrated2Zulip
Photon Commissioning Team
XRO
DOC
HED Migrated2Zulip
WP74
MID
PSPO_XO
XO_Operation
SQS Migrated2Zulip
FXE Migrated2Zulip
DET
Sample Environment
SPB Migrated2Zulip
OPERATION-2025
OPERATION-2024
OPERATION-2023
OPERATION-2022
OPERATION-2021
OPERATION-2020
OPERATION-2019
SASE1-2018
SASE1-2017
Experiment data
OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn
Not logged in
List
|
Message ID:
18
Entry time:
10 Feb 2024, 00:12
Author:
Tokushi Sato
Group:
SPB
Subject:
Shift summary
X-ray delivery
1-350 at 12.45 keV, 1700 uJ. Stable delivery
Optical laser delivery
None
Achievements/Observations
collected about over 15 hours of quality data from different samples (thanks to FXE) with extremely stable jet.
Issues
Issue with data processing in the morning and from the day before due to DAQ; only half of the detector data were calibrated; this was solved by switching to the old DAQ
ELOG V3.1.4-7c3fd00